OpenStudy (anonymous):

In triangle ABC, A = 62 degrees, B = 56 degrees, and c = 28. What is b?

OpenStudy (cggurumanjunath):

USE COSINE FORMULA !

OpenStudy (jdoe0001):

\(\bf \textit{Law of Cosines}\\ c^2= a^2+b^2-2abcos(C)\implies c=\sqrt{a^2+b^2-2abcos(C)}\)

OpenStudy (cggurumanjunath):

CAN U DRAW THE FIGURE ?